Supply and Demand (8 points)

Consider the market for silver, where quantity is in ounces, and price in dollars.

Price (P)

$27

$24

$21

$18

$15

$12

$6

$3

$0

Quantity Demanded (QD)

0

2

4

6

8

10

12

14

16

Quantity Supplied (QS)

16

14

12

10

8

6

4

2

0

a) What are the equilibrium price and quantity? Why?

b) Assume that the Government puts in place a price ceiling (maximum) of $6 per ounce. What will be the new Qs, the new Qd, the actual amount bought/sold, and the shortage/surplus (if any)? Why?

c) Assume that the government imposes a price floor (minimum) of $ 21 per ounce in this market. What will be the new Qs, the new Qd, the actual amount bought/sold, and the shortage/surplus (if any)? Why?

d) Calculate the total revenue received by the seller from actual trade (defined as actual price multiplied by actual quantity) in this market in the original equilibrium, and under the price controls described in parts b and c above. Explain.
Answer #1

Ans

Equilbrium price is 15 and quantity 8 because here there is no excess supply or demand because supply=demand 2 Qs=4 and Qd=12.There is shortage=12-4=8 because demand is greater than supply

3 Qs=14 and Qd=2.There is excess supply=14-2=12 because demand is less than supply

4 revenue from actual trade=15(8)=120

Revenue from b=6(4)=24

Revenue from c=2(21)=42